Prove that tanA + 2 tan2A + 4 tan4A + 8 cot8A = cotA

Prove that 


      tanA + 2 tan2A + 4 tan4A + 8 cot8A = cotA

Grade:11

3 Answers

pavani suhani
12 years ago

cotA-tanA=2cot2A

            cotA-cotA+tanA+2tan2A+4tan4A+8cot8A

            =cotA-2cot2A+2tan2A+4tan4A+8cot8A

            =cotA-4cot4A+4tan4A+8cot8A

            =cotA-8cot8A+8cot8A

             =cotA

Ashwin Muralidharan IIT Madras
12 years ago

Hi Ankit,

Best Regards,

Ashwin (IIT Madras).

ankit agrawal
12 years ago

tan A + 2 tan2A + 4 tan4A + 8 cot8A = cotA